# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v2

@@ 7.3 Random Forests: explanation multiplicity in ensemble models
Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class. Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling. Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more expressive and widely deployed ensemble architectures.
@@ 7.3.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Fig. 7. Each train–test split aggregates approximately Random Forest instances, reflecting stochasticity arising from bootstrap resampling, feature subsampling, and hyperparameter variation. This large-scale sampling allows the structure of explanation space to be examined at high resolution.
Figure 6 visualises the universal explanation manifold obtained by stacking explanation vectors across multiple train–test splits. In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into multiple dense and well-separated regions. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting. As in the linear case, the global average attribution vector—computed by averaging SHAP values across all runs—lies outside the high-density regions of the manifold. This indicates that aggregating explanations across stochastic training runs again produces a representation that does not correspond to a typical deployed model instance.
Importantly, predictive accuracy varies smoothly across the explanation manifold and does not align with any single cluster or region. Models drawn from different explanation clusters achieve comparable test accuracy, with no sharp performance boundary separating the modes. This demonstrates that, in Random Forests, explanation multiplicity is not driven by accuracy trade-offs or poorly performing solutions, but instead reflects the coexistence of distinct explanatory mechanisms within the same performance regime.
Figure 7 shows the explanation landscape for a single representative split (Split 104). Even within a fixed train–test realisation, explanation vectors again separate into two high-density regions. The first principal component captures the majority of variance in explanation space, while the second component reflects residual spread within each mode. As in the universal manifold, predictive accuracy varies continuously across both regions, indicating that explanation multiplicity arises intrinsically within a single split and does not depend on aggregation across data resamplings or extreme hyperparameter configurations.
@@ 8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (Lipton (2018); Rudin (2019)). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
@@
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others (2020); Selbst and Barocas (2018)).
@@
EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or optimisation randomness. Instead, it makes this uncertainty visible. The framework is deliberately agnostic to causal origin and instead focuses on empirical detectability.
@@
The multiplicity observed cannot be attributed to instability or noise in SHAP itself. SHAP values are deterministic for any fixed model. EvoXplain examines attribution profiles across independent model trainings, rather than variability within a single explanation. This distinction is critical: EvoXplain characterises variability across trained models, not variability within an attribution method. Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained models rather than stochasticity in the explainer itself. As such, the resulting clusters reflect real differences in learned mechanisms. For the Breast Cancer benchmark, the centroids of the discovered basins reinforce established diagnostic attributes such as perimeter, radius, area, concavity, and associated morphological descriptors used in clinical tumour assessment. The emergence of distinct, clinically meaningful feature profiles across basins would not be expected if clustering were driven by random noise.
@@
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission (2023); Rudin (2019)).
@@
This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al. (2021); Habli et al. (2025)), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone. Within such frameworks, explanation stability is not merely an interpretability preference but a form of verification evidence. EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al. (2025)).
@@
This is particularly important because explanations have immediate consequences in certain domains. In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification Umbach et al. (2023). In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims Agrawal and Jain (2023). Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining. In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
In our experiments, we focus on tabular learning, where repeated training is manageable. However, retraining large language models (LLMs) presents a more difficult challenge. We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings. LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially. Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question. Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (Olah et al. (2020)). EvoXplain provides a conceptual starting point. The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (Lakshminarayanan et al. (2017)).
Beyond the empirical evidence, mechanistic multiplicity underscores the need to evaluate reliability and interpretability as properties that cannot be assumed. Instead, they must be empirically established under realistic training variability.
Why Mechanistic Multiplicity Matters
Mechanistic multiplicity primarily raises questions of reliability, rather than philosophy. It concerns whether a modelling pipeline produces consistent reasoning under retraining, not whether a single explanation appears plausible. When predictive accuracy decays, retraining is standard practice, and the modelling process may generate different reasoning outputs, with no guarantee of stable reasoning over time. Reliability must therefore concern the process rather than a particular model instance. In high-stakes contexts, explanations justify actions. Even at equal predictive accuracy, models can diverge in the actions, risk factors, and scientific hypotheses they imply, settling into different basins of explanation. This divergence is not resolved by appeals to sparsity, simplicity, or accuracy alone. When explanations inform decisions rather than merely describe them, instability in the underlying mechanism becomes a substantive safety and interpretability issue. Appeals to simplicity do not resolve multiplicity. In the presence of competing explanation basins with equivalent sparsity, complexity, and predictive accuracy, Occam’s razor cannot favour one over another. What remains is a form of empirical underdetermination. In many cases, competing explanation basins are comparable not only in accuracy, but also in sparsity and apparent simplicity, leaving no principled criterion for selection based solely on model-internal properties. In such cases, the presence of multiple stable explanatory modes evidences a profound underdetermination, which EvoXplain is designed to identify and document.
There are boundaries to this work. We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (Olah et al. (2020)). EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes Selbst and Barocas (2018). We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling. These limitations define the scope of the present study rather than its validity. The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
Despite these limitations, EvoXplain opens new research questions. Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (Marx et al. (2020))? Is it possible to prefer certain basins for safety, fairness, or generalisability (Rudin (2019))? More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure. This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.
To summarise, EvoXplain treats interpretability as a property of model behaviour under retraining. By focusing on mechanistic multiplicity rather than assuming a single explanatory mode, EvoXplain provides tools to uncover and characterise multiplicity when it arises. This, in turn, raises methodological, epistemic, and governance challenges. If models admit multiple internally consistent mechanistic explanations, then explanation selection, validation, and governance require explicit criteria. If models embody a multiplicity of mechanistic stories, how should they be assessed, described, or governed? While EvoXplain does not resolve these questions, it makes them explicit and positions explanation stability as a concrete target for future research. In practice, EvoXplain enables practitioners and auditors to detect when single-model explanations are insufficient, without prescribing which explanation should be preferred.
@@ 9 Conclusion
Explanations are often treated as fixed model properties—examined once and assumed to reflect one underlying mechanism. EvoXplain argues the opposite, and instead quantifies this assumed stability. Across two benchmarks and two model families, models achieved high predictive performance. However, their explanations frequently diverged into multiple distinct modes, even within the same model class and data split. The extent and structure of this multiplicity varied by dataset—for example, explanation basins were more evenly balanced on the Breast Cancer dataset and more asymmetric on COMPAS—but in all cases, high predictive accuracy did not imply a single stable mechanism. This demonstrates that explanation stability cannot be inferred from predictive performance, model class, or perceived simplicity alone.
EvoXplain offers a diagnostic view rather than a decision rule. It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge. It does not dictate which local minima should be prioritized; that is a decision to be made from the domain knowledge, validation or regulatory governance. It is a valuable tool to make multiplicity non-abstract and measurable, and to encourage interpretability practices that account for stability. In this sense, EvoXplain complements existing interpretability methods by exposing when single-model explanations are insufficient.
@@
Interpretability can no longer centre on a single trained model. The assumed single explanation is viewed as a distribution of explanations, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement. EvoXplain pushes the field in this direction by treating explanation stability as a primary diagnostic of model behaviour. By making explanation multiplicity observable and quantifiable, EvoXplain provides a foundation for future interpretability, auditing, and governance frameworks that operate at the level of model populations rather than individual fits.

### v3

@@ 7.3 Random Forests: explanation multiplicity in ensemble models
We next evaluate EvoXplain on Random Forests, a model class that introduces nonlinearity, ensemble averaging, and intrinsic stochasticity through bootstrap resampling and feature subsampling. This setting allows us to test whether explanation multiplicity persists when predictive variance is strongly suppressed by ensembling, using five independent train–test splits.
@@ 7.3.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Fig. 7. We train a population of Random Forest instances per train–test split by sampling stochastic sources intrinsic to the algorithm (bootstrap resampling and feature subsampling).
Figure 6 visualises the universal explanation manifold obtained by stacking explanation vectors across multiple train–test splits. In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into two well-separated regions in explanation space, with partial overlap at their boundaries. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting. As in the linear case, the global average attribution vector—computed by averaging SHAP values across all instances—lies outside the high-density regions of the manifold. This indicates that aggregating explanations across admissible training pipeline instantiations produces a representation that does not correspond to any high-density region of the explanation distribution.
Importantly, predictive accuracy does not align with any single basin or region. Instances drawn from different explanation basins achieve comparable test accuracy, with no sharp performance boundary separating the modes. This demonstrates that, in Random Forests, explanation multiplicity is not driven by accuracy trade-offs or poorly performing solutions, but instead reflects the coexistence of distinct explanatory mechanisms within the same performance regime.
Figure 7 shows the explanation landscape for a single representative split (Split 104). Even within a fixed train–test realisation, explanation vectors again form two statistically separable regions in the explanation space. The first principal component captures the majority of the variance in explanation space, while the second component reflects the residual spread within each mode. As in the universal manifold, predictive accuracy remains comparable across these regions, indicating that explanation multiplicity is already present within a single train–test split and is not an artefact of aggregating explanations across multiple data resamplings.
@@ 8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining and model selection procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when the training pipeline admits multiple admissible configurations (here instantiated via regularisation variation), despite high and stable predictive accuracy (Lipton, 2018; Rudin, 2019). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
@@
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others, 2020; Selbst and Barocas, 2018).
@@
EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or the non-identifiability of admissible training configurations within a pipeline. Instead, it makes this uncertainty visible. The framework is deliberately agnostic to causal origin and instead focuses on empirical detectability.
Deterministic models can still exhibit pipeline-induced multiplicity.
A potential source of confusion is that regularised Logistic Regression with a deterministic solver is uniquely determined by a fixed dataset, preprocessing, and a fixed regularisation strength . EvoXplain does not assume that a deployed system samples only from this single point. In practice, retraining and model updates include upstream degrees of freedom such as hyperparameter selection, data refresh, and preprocessing changes, which together induce a distribution over admissible model instantiations (D’Amour et al., 2022; Breiman, 2001; Marx et al., 2020). Consistent with this framing, an orthogonal synthetic control confirms that with fixed explanations collapse to , whereas allowing to vary reintroduces multimodality, indicating that multiplicity arises from the instantiation pipeline rather than solver randomness.
Logistic Regression is the minimal test case.
It is important to emphasise that -regularised Logistic Regression represents the best-case scenario for explanation stability: the loss is convex, the solver is deterministic, and the model is fully identifiable once the regularisation strength is fixed. If explanation multiplicity arises even in this setting when the deployment pipeline admits multiple admissible instantiations, then the phenomenon is necessarily more pronounced in model classes with intrinsic stochasticity or non-convexity. Random Forests introduce bootstrap and feature-subsampling randomness; Elastic Net regularisation admits multiple sparse solutions under correlated features; and deep neural networks exhibit vast equivalence classes of parameterisations with similar predictive behaviour. EvoXplain therefore establishes a lower bound on explanation multiplicity: if a pipeline fails to identify a unique explanatory regime in the simplest convex case, it cannot be expected to do so in more flexible or stochastic architectures. Importantly, the presence of multiple explanation basins at comparable hyperparameter values—for example, basins with nearly identical mean regularisation strength in Logistic Regression (Table 1)—rules out an interpretation of EvoXplain as merely measuring continuous attribution drift, and instead points to discrete explanation non-identifiability within admissible training pipelines.
@@
The multiplicity observed cannot be attributed to instability or noise in SHAP itself. SHAP values are deterministic for any fixed model. EvoXplain examines attribution profiles across independently trained model instances produced by the same pipeline, rather than variability within a single explanation. This distinction is critical: EvoXplain characterises variability across trained models, not variability within an attribution method. Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained model instances rather than stochasticity in the explainer itself. As such, the resulting basins reflect real differences in learned parameterisations and their induced attribution behaviour. For the Breast Cancer benchmark, the centroids of the discovered basins reinforce established diagnostic attributes such as perimeter, radius, area, concavity, and associated morphological descriptors used in clinical tumour assessment. The emergence of distinct, clinically meaningful feature profiles across basins would not be expected if clustering were driven by random noise.
@@
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission, 2023; Rudin, 2019).
@@
This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al., 2021; Habli et al., 2025), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone. Within such frameworks, explanation stability is not merely an interpretability preference but a form of verification evidence. EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining and model selection procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al., 2025).
@@ 9 Conclusion
Explanations are often treated as fixed model properties—examined once and assumed to reflect one underlying mechanism. EvoXplain argues the opposite, and instead quantifies this assumed stability under realistic retraining and model selection procedures. Across two benchmarks and two model families, models achieved high predictive performance. However, their explanations frequently diverged into multiple distinct modes whenever the training pipeline admitted more than one near-equivalent solution, even within the same model class and data split. The extent and structure of this multiplicity varied by dataset—for example, explanation basins were more evenly balanced on the Breast Cancer dataset and more asymmetric on COMPAS—but in all cases, high predictive accuracy did not imply a single stable mechanism. This demonstrates that explanation stability cannot be inferred from predictive performance, model class, or perceived simplicity alone.
EvoXplain offers a diagnostic view rather than a decision rule. It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge under admissible retraining. It does not dictate which admissible solution should be prioritized; that is a decision to be made from domain knowledge, validation or regulatory governance. It is a valuable tool to make multiplicity non-abstract and measurable, and to encourage interpretability practices that account for stability. In this sense, EvoXplain complements existing interpretability methods by exposing when single-model explanations are insufficient.
@@
Interpretability can no longer centre on a single trained model. The assumed single explanation should be viewed as a distribution of explanations induced by the training pipeline, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement. EvoXplain pushes the field in this direction by treating explanation stability as a primary diagnostic of model behaviour. By making explanation multiplicity observable and quantifiable, EvoXplain provides a foundation for future interpretability, auditing, and governance frameworks that operate at the level of model populations rather than individual fits.
Appendix A Supplementary Materials and Reproducibility
All supplementary results, including robustness checks, synthetic controls, and full clustering outputs, are provided in the public code repository.111https://github.com/bensmailchama-boop/EvoXplain/
The repository contains result folders generated directly by the EvoXplain pipeline, including: (i) per-split clustering summaries, (ii) silhouette and entropy values, (iii) PCA visualisations with basin overlays, and (iv) metadata files specifying experimental configurations.